1. One card is choesen from a standard deck.
a. what is the probability of less than 7 (Ace high)? i really need help for this question.

Answer by stanbon(57203)

One card is choesen from a standard deck.
a. what is the probability of less than 7
There are 5 cards less than 7 in each of 4 suits = 28 cards
---
P(less than 7) = 28/52 = 19/26
=============
Cheers,
Stan H.